OpenStudy (pottersheep):

Find a quadratic equation with the roots 3 + 4i and 3 - 4i. HELP PLEASE!

OpenStudy (amistre64):

when given roots; just reconstruct the poly with them

OpenStudy (amistre64):

x=(a,b) (x-a)(x-b)=0

OpenStudy (ash2326):

if we are given roots as a and b , then the equation is given as (x-a)(x-b)=0 here the roots are 3+4i and 3 -4i (x-(3+4i))(x-(3-4i)) x^2-(3+4i+3-4i)x-(9-16)=0 x^2-6x-7=0

OpenStudy (amistre64):

if anything:\[x=\frac{-b+\sqrt{b^2-4ac}}{2a}\]

OpenStudy (amistre64):

3 = -b/2a -3/2 = b/a 2x^2 -3x + c = 0 is a sort of way to see it

OpenStudy (amistre64):

except for I mistyped it lol

OpenStudy (amistre64):

-6/1 = b/a x^2 -6x + c = 0

OpenStudy (pottersheep):

but my answers say x^2 + 6x + 25 :s

OpenStudy (anonymous):

The simplest "Vieta's Formula" is x^2 - (sum)x + (product) in this case we get x^2 - 6x + 25